Possible Outcomes when die is rolled
dot showed by die , possible outcome in single roll
thus , formula for the probability distribution of the random variable x will be
P(X=x) = x/6Cx Where as x = 1
A coin is tossed 2 times what are the number of possible outcomes?
There are 4 possible outcomes.
There are 2 outcomes (heads or tails) on the first toss and 2 on the second toss.
The possibilities are HH, TT, HT and TH.
C(40,2)c(12,0)/c(52,2) = 0.588235294
What are the odds of getting 6 tails on 6 coin tosses?
If the coin is fair, then
The chance of tails on the 1st toss is 50% (1/2).
The chance of tails on the 2nd toss is 50% (1/2).
The chance of tails on the 3rd toss is 50% (1/2).
The chance of tails on the 4th toss is 50% (1/2).
The chance of tails on the 5th toss is 50% (1/2).
The chance of tails on the 6th toss is 50% (1/2).
The probability of getting all of them is (1/2) x (1/2) x (1/2) x (1/2) x (1/2) x (1/2) = 1/64 = 1.5625 % .
The odds are 63 to 1 against.
7/(4+7+5) = 7/16 = 43.75%
The mean is 12 and each observation is 8 units away from 12.
1 out of 15
Probab. = Prob. of red x Prob. of blue
Probab. = (3/10)x(2/9) = 5/90 = 1/15
You roll two number cubes what is the probability that you roll a 1 first and a 2 second?
Since there are 6 sides on every die that are equally likely to be rolled, the probability of rolling any given side once is exactly 1/6. The 2 events or the first and second dice roll are independent (the outcome of one does not influence the other) so to find the probability of both occurring you just multiply the probability of each event. Since each event has a 1/6 probability of occurring as stated before, The entire event has a probability of 1/6*1/6 or 1/36, which is approximately 2.78%.
What is the difference standard error of mean and sampling error?
The standard error of the mean and sampling error are two similar but still very different things.
In order to find some statistical information about a group that is extremely large, you are often only able to look into a small group called a sample. In order to gain some insight into the reliability of your sample, you have to look at its standard deviation. Standard deviation in general tells you spread out or variable your data is. If you have a low standard deviation, that means your data is very close together with little variability. The standard deviation of the mean is calculated by dividing the standard deviation of the sample by the square root of the number of things in the sample. What this essentially tells you is how certain are that your sample accurately describes the entire group. A low standard error of the mean implies a very high accuracy.
While the standard error of the mean just gives a sense for how far you are away from a true value, the sampling error gives you the exact value of the error by subtracting the value calculated for the sample from the value for the entire group. However, since it is often hard to find a value for an entire large group, this exact calculation is often impossible, while the standard error of the mean can always be found.
What is a space diagram in maths probability?
A space diagram is commonly used in mathematics. It is a
table which represents a range of work to mostly do with probability!
Hope it helps
z score = (400-500)/80=-1.25
probability under curve to right of -1.25 = .8944
i think the answer is 9
How many ways can you arrange 4 objects?
The first object in the row can be any one of the 4 objects. For each of those . . .
The second one in the row can be any one of the remaining 3 objects. For each of those . . .
The third one in the row can be either of the remaining 2 objects.
The total number of different arrangements is (4 x 3 x 2) = 24 ways.
Suppose you toss both plastic coins 1000 times instead of 100 what would this larger sample allow?
The Larger sample allows for more precise results.
How many possible outcomes can you get by tossing 5 coins?
There are 25 or 32 possible outcomes can you get by tossing 5 coins.
Any 2 of 12 is (12 x 11)/2 = 66. Any 2 from 6 is (6 x 5)/2 = 15 so the probability is 15 out of 66 ie 0.227 or 22.7%
The number is 8!/[3!(8-3)!] where n! = 1*2*3*...*n
ie 8*7*6/(3*2*1) = 56
What does it mean if the standard deviation is less than the mean?
If the mean is less than or equal to zero, it means there has been a serious calculation error.
If the mean is greater than zero and the distribution is Gaussian (standard normal), it means that there is an 84.1% chance that the value of a randomly variable will be positive.
.99
How does sample size affect validity of results in a research?
A large sample reduces the variability of the estimate. The extent to which variability is reduced depends on the quality of the sample, what variable is being estimated and the underlying distribution for that variable.